OpenStudy (anonymous):

Find the sum of the geometric sequence –3, 15, –75, 375, … when there are 8 terms and select the correct answer below.

OpenStudy (anonymous):

–195,312 195,312<-----my answer 292,968 –292,968

OpenStudy (campbell_st):

good choice.

OpenStudy (anonymous):

a=3 \[r=\frac{ 15 }{ 3 }=-5\] \[S _{n}=\frac{ a \left( r^n-1 \right) }{ r-1 }\] \[S _{8}=\frac{ -3\left\{ \left( -5 \right)^8-1 \right\} }{ -5-1 }\] ?